ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 263, 2024

Araújo, J., P. J. Cameron, C. Casolo, F. Matucci and C. Quadrelli Integrals of groups. II

49

Bor, G. and T. Willse Chains of path geometries on surfaces: theory and examples

587

Bouabdillah, O. and C. Le Merdy Polygonal functional calculus for operators with finite peripheral spectrum

517

Bowen, L. and Y. F. Lin Entropy for actions of free groups under bounded orbit-equivalence

809

Cameron, P. J. See J. Araújo Casolo, C. See J. Araújo Cossu, L. and S. Tringali Abstract factorization theorems with

applications to idempotent factorizations

349

Cui, W., L. Luo and W. Wang Cells in affine q-Schur algebras  129
Das, S. Galois covers of singular curves in positive characteristics  397

d’Elbée, C., I. Kaplan and L. Neuhauser On algebraically closed fields with a distinguished subfield

169

Fox, J., X. He and Y. Wigderson Ramsey numbers of sparse digraphs

1

Grbac, N. and H. Grobner Some unexpected phenomena in the Franke filtration of the space of automorphic forms of the general linear group

301

Grobner, H. See N. Grbac Guo, Y. and S. Peng Maximum principles and direct methods for

tempered fractional operators

693

Guzmán, O. and S. Todorcevic The P -ideal dichotomy, Martin’s axiom and entangled sets

909

Halevi, Y. and A. Hasson Corrigendum to “Strongly dependent ordered abelian groups and henselian fields”

433

Hasson, A. See Y. Halevi He, X. See J. Fox Higueras-Montaño, L. F. and N. Jonard-Pérez A topological

insight into the polar involution of convex sets

437

Jonard-Pérez, N. See L. F. Higueras-Montaño Kaplan, I. See C. d’Elbée Ko, H., V. Mazorchuk and R. Mr-den Join operation for the

Bruhat order and Verma modules

627

Kus, D. and R. Venkatesh Simplified presentations and embeddings of Demazure modules

475

Lin, Y. F. See L. Bowen Luo, L. See W. Cui Marshall-Maldonado, J. Modulus of continuity for spectral mea-

sures of suspension flows over Salem type substitutions

965

Matucci, F. See J. Araújo Mazorchuk, V. See H. Ko Mr-den, R. See H. Ko Le Merdy, C. See O. Bouabdillah Neuhauser, L. See C. d’Elbée Peng, S. See Y. Guo Quadrelli, C. See J. Araújo Shani, A. Strong ergodicity around countable products of countable

equivalence relations

871

Steinerberger, S. Effective bounds for the decay of Schrödinger eigenfunctions and Agmon bubbles

781

Todorcevic, S. See O. Guzmán

Towsner, H. and K. Yokoyama Erdős–Moser and IΣ_{2}  843

Tringali, S. See L. Cossu Vassiliev, V. A. Complements of discriminants of simple real

function singularities

553

Venkatesh, R. See D. Kus Wang, W. See W. Cui

Weilacher, F. Borel edge colorings for finite-dimensional groups  737

Wigderson, Y. See J. Fox Willse, T. See G. Bor Winkler, M. Complete infinite-time mass aggregation in a quasilinear

Keller–Segel system

93

Yokoyama, K. See H. Towsner Yu, Z. Group-theoretical property of some integral non-degenerate

fusion categories

501

Zavyalov, B. Altered local uniformization of rigid-analytic spaces  213